# Invisible Labor in Open Source Software Ecosystems

### v1

@@ 5.2. Limitations & Future Work
Our work has several limitations. First of all, survey instruments tend to prove more reliable in the form of validated scales. Questions remain about how feasible creating a scale would be for something so difficult to trace which motivated the exploratory nature of this study. That said, existing information visibility scales (ter Hoeven et al., 2021) may guide the labor visibility portion, leaving only need for compensation measures. Next, our sample sizes are too small to make these findings conclusive about all of OSS. In particular, we did not see differences between communities as we might have expected; breaking down our responses by artifacts and use cases did not show statistical differences between category means on any of the quantitative questions (0 of 16, 0%). Sampling OSS ecosystems broadly as we did here may prove less helpful going forward than conducting targeted studies to benefit particular communities, though. So future efforts might focus on specific artifacts, use cases, geographic regions, or programming language communities. Third, our sample of convenience via social media prevented us from engaging with those who use social media less frequently meaning our sample may contain biases toward social media users and more socially-engaged ecosystems.
@@
Perhaps the greatest need arising from this work is a granular understanding of how specific artifacts facilitate the existence of labor data, its sharing, accessibility, and participant compensation. For example, we saw that certain attribution mediums corresponded to greater credit satisfaction. Further research might confirm whether appropriately selecting credit mediums improves accessibility of attribution and, in turn, improves worker satisfaction. These artifacts also shape worker perceptions of visibility, so researchers and practitioners likely need to pay careful attention to the interplay between accessibility and perceptions of openness. Labor will not become fully visible without ensuring consistent support for each factor that limits labor visibility and compensation, so articulation work and infrastructuring activities would benefit from renewed attention.
@@
And of course, one study of invisibility will not make all labor visible, so future work might triangulate our findings within OSS through other qualitative and quantitative methods. Studies of invisible labor need not limit themselves to OSS contexts either as invisible labor happens everywhere, in computing and more broadly. For CSCW, this article provides further evidence in support of what Toxtli & colleagues (Toxtli et al., 2021) recently found in crowd work. We hope others will similarly build on our work in OSS and other domains.
@@ 6. Conclusion
Much of the work that happens in OSS ecosystems is not visible, not compensated, or both. Such invisible labor makes it difficult for individuals to advance their careers and for organizations to understand processes that enable their organizations to run. In this study, we measured invisible labor in OSS ecosystems by estimating the visibility and compensation rates of OSS activities. We found that approximately half of OSS labor does not receive credit, a low-cost form of attribution that often precedes other forms of compensation, and that OSS activities are more likely to be nonvisible or partially visible than to be fully visible. In addition, we demonstrated that anchoring labor to the concept of visibility tends to increase visibility frequency perceptions and decrease credit importance assessments. These findings tell us that compensating all those who work in OSS ecosystems will require careful attention to documenting labor activities, as well as sharing and making documentation accessible.
@@ Acknowledgements.
Thanks to Jean-Gabriel Young and Melania Lavric for their suggestions throughout. The authors were partially supported by Google under the Open Source Complex Ecosystems and Networks project. J.M. was partially supported by the Sloan Foundation through the Vermont Research Open Source Program Office, and by the NSF award BCS-2332598.

### v2

@@ 6.2. Limitations & Future Work
Our work has several limitations. First of all, survey instruments tend to prove more reliable in the form of validated scales. Questions remain about how feasible creating a scale would be for something so difficult to trace which motivated the exploratory nature of this study. That said, existing information visibility scales (ter Hoeven et al., 2021) may guide the labor visibility portion, leaving only need for compensation measures. Next, our sample sizes are too small to make these findings conclusive about all of OSS. In particular, we did not see differences between communities as we might have expected; breaking down our responses by artifacts and use cases did not show statistical differences between category means on any of the quantitative questions (0 of 16, 0%). Sampling OSS ecosystems broadly as we did here may prove less helpful going forward than conducting targeted studies to benefit particular communities or study specific ecosystem characteristics, though. We were also limited in the questions we could ask participants because our survey spanned multiple ethnics regulatory jurisdictions. So future efforts might focus on specific artifacts, use cases, geographic regions, programming language communities, sizes of projects and ecosystems, volunteer statuses, corporate participation, years of participation, membership status, regulatory jurisdictions, and many other variables that were difficult to measure through an anonymous, self-reported survey. Third, our sample of convenience via social media prevented us from engaging with those who use social media less frequently meaning our sample may contain biases toward social media users and more socially-engaged ecosystems.
@@
Perhaps the greatest need arising from this work is a granular understanding of how technology affordances, co-evolving technologies and goals (Leonardi, 2011; Gibson et al., 2022), shape visibility and invisibility in OSS ecosystems. How do specific artifacts facilitate the existence of labor data, its sharing, accessibility, and participant compensation? And how do participants’ motivations shift as a product of the capabilities of the resulting artifact? For example, we saw that certain attribution mediums corresponded to greater credit satisfaction. This was further supported by our qualitative analysis wherein participants preferred more or less automated and personalized attribution depending on their motivations. Further research might examine how interaction between credit mediums and motivations shape worker satisfaction (though other factors may influence satisfaction as our findings only show a correlation). These artifacts also shape worker perceptions of visibility, so researchers and practitioners likely need to pay careful attention to the interplay between accessibility and perceptions of openness. Labor will not become visible (to the desired level of stakeholders in each ecosystem) without ensuring consistent support for each factor that influences labor visibility and compensation. So, articulation work and infrastructuring activities would benefit from renewed attention, as would economic implications of whether or not invisible labor receives compensation.
@@
Future work might also triangulate our findings within OSS through other qualitative and quantitative methods. We especially invite researchers to explore other factors that may influence invisible labor including technology affordances (Leonardi, 2011), gender (Frluckaj et al., 2022; Berdahl et al., 2018), governance (Shah, 2006; Frey et al., 2019), and ecosystem size (Geiger et al., 2021), each of which may have some effect. Studies of invisible labor need not limit themselves to OSS contexts either as invisible labor happens everywhere, in computing and more broadly. For CSCW, this article provides further evidence in support of what Toxtli & colleagues (Toxtli et al., 2021) recently found in crowd work. We hope others will similarly build on our work in OSS and other domains.
@@ 7. Conclusion
Much of the work that happens in OSS ecosystems is not visible, not compensated, or both. Such invisible labor makes it difficult for individuals to advance their careers, delegitimizies “background work,” and limits organizational understanding of crucial processes. In this study, we measured invisible labor in OSS ecosystems by estimating the visibility and compensation rates of OSS activities. We found that approximately half of OSS labor does not receive credit, a low-cost form of attribution that often precedes other forms of compensation, and that conflicting motivations amidst attribution practices may be partially responsible. In addition, we demonstrated that anchoring labor to the concept of visibility may lead participants to overestimate the visibility of their work and decrease its importance to them. Designing attribution and compensation systems with diverse stakeholder groups — while assuming that most OSS work is not visible or compensated — may therefore help OSS ecosystems better satisfy the varied motivations of their participants.
@@ Acknowledgements.
Thanks to Jean-Gabriel Young, Melania Lavric, and the anonymous reviewers for their suggestions throughout. The authors were partially supported by Google under the Open Source Complex Ecosystems and Networks project. J.M. was partially supported by the Sloan Foundation through the Vermont Research Open Source Program Office, and by the NSF award BCS-2332598. M.Z.T. was partially supported by the Northeastern University Future Faculty Postdoctoral Fellowship Program.